Describe the transformations on the following graph of f(x)=log(x). State the placement of the vertical asymptote and x-intercept after the transformation. For example, vertical shift up 2 or reflected about the x-axis are descriptions.
a) g(x)=log(x-3)
b) g(x)=log(-x)

Answer by stanbon(75887)

Describe the transformations on the following graph of f(x)=log(x). State the placement of the vertical asymptote and x-intercept after the transformation. For example, vertical shift up 2 or reflected about the x-axis are descriptions.
Equation(s) for the Vertical Asymptote(s):
x-intercept in (x, y) form
a) g(x)=log(x-3)
Ans: shifted 3 units to the right.
vertical asymptote: x = 3
x-intercept: Let g(x)=0; log(x-3)=0; x-3 = 10^0; x = 4
---------------------------------------
b) g(x)=log(-x)
Ans: reflected in the y-axis.
vertical asymptote: y = 0
x-intercept: Let g(x) = 0, then log(-x)=0 ; -x=10^0 = 1; x = -1
=================================
Cheers,
Stan H.